### v2

# Logistic-Beta Processes for Dependent Random Probabilities with Beta Marginals

Abstract
The beta distribution serves as a canonical tool for modelling probabilities in statistics and machine learning. However, there is limited work on flexible and computationally convenient stochastic process extensions for modelling dependent random probabilities. We propose a novel stochastic process called the logistic-beta process, whose logistic transformation yields a stochastic process with common beta marginals. Logistic-beta processes can model dependence on both discrete and continuous domains, such as space or time, and have a flexible dependence structure through correlation kernels. Moreover, its normal variance-mean mixture representation leads to effective posterior inference algorithms. We illustrate the benefits through nonparametric binary regression and conditional density estimation examples, both in simulation studies and in a pregnancy outcome application.
Keywords: Bayesian nonparametrics; Dependent Dirichlet process; Multivariate beta distribution; Nonparametric binary regression; Pólya distribution.
1 Introduction
The beta distribution is widely used for modelling random probabilities owing to its flexibility, natural interpretation of parameters, and conjugacy in Bayesian inference. Beta distributions play a pivotal role in Bayesian nonparametric modelling (Hjort et al.,, 2010; Müller et al.,, 2015; Ghosal and van der Vaart,, 2017). For instance, a broad class of random probability measures have stick-breaking representations in terms of beta random variables, including the popular Dirichlet processes and Pitman-Yor processes (Ferguson,, 1973; Sethuraman,, 1994; Perman et al.,, 1992; Pitman and Yor,, 1997; Ishwaran and James,, 2001). Other notable models that build upon the beta distribution include Pólya trees (Ferguson,, 1974) for nonparametric density estimation and beta processes (Hjort,, 1990), which were originally designed as priors for the cumulative hazard rate but have been used broadly in machine learning due to their link with latent feature models, such as the Indian buffet process (Griffiths and Ghahramani,, 2011).
Stochastic processes with beta marginals play a pivotal role in extending such models to handle non-exchangeable and heterogeneous data sources, while maintaining marginal properties; examples include dependent Dirichlet processes (MacEachern,, 1999; Quintana et al.,, 2022), dependent Pólya trees (Trippa et al.,, 2011) and dependent Indian buffet processes (Perrone et al.,, 2017). This article is motivated by the need to define flexible and computationally convenient classes of beta processes to substantially improve over current approaches, such as copula-based formulations (MacEachern,, 2000; Arbel et al.,, 2016; De Iorio et al.,, 2023). Copulas are highly flexible but lead to challenging posterior computation, particularly in complex settings. Trippa et al., (2011) and Bassetti et al., (2014) instead extend bivariate beta distributions (Olkin and Liu,, 2003; Nadarajah and Kotz,, 2005), but this can only induce positive correlation. Other strategies rely on latent binomial random variables (Nieto-Barajas et al.,, 2012) and transformation of Gaussian autoregressive processes (DeYoreo and Kottas,, 2018), but focused narrowly on discrete-time dependence. Finally, constructions based on covariate-dependent ordering (Griffin and Steel,, 2006), local sharing of random components (Chung and Dunson,, 2011), and Wright-Fisher diffusion (Perrone et al.,, 2017) are complex and limited to specialized cases.
We propose a novel logistic-beta process, whose logistic transformation is a stochastic process with common beta marginals. The logistic-beta process builds upon the normal variance-mean mixture representation of a univariate logistic-beta distribution (Barndorff-Nielsen et al.,, 1982), and extends its hierarchical formulation to a stochastic process through correlation kernels in a related manner to a Gaussian process. As such, the proposed process can be flexibly defined on either discrete or continuous domains, and the implied dependence ranges from a perfect correlation to possibly negative correlations. Adopting the logistic-beta process prior for logit-transformed latent random probabilities provides a conditionally conjugate posterior sampling scheme that can directly exploit Pólya-Gamma data augmentation strategies (Polson et al.,, 2013).
We provide two modeling examples utilizing the logistic-beta process: nonparametric regression for dependent binary data and conditional density estimation via a dependent Dirichlet process mixture with covariate-dependent weights and atoms. There are few methods for covariate-dependent weights in the dependent Dirichlet process due to computational challenges (Quintana et al.,, 2022; Wade et al.,, 2023). We illustrate how the proposed formulation leads to efficient posterior inference algorithms. Although there are alternative Bayesian nonparametric models for conditional densities (Dunson and Park,, 2008; Ren et al.,, 2011; Rigon and Durante,, 2021), maintaining a marginal Dirichlet process structure has advantages in terms of interpretability, ease of modelling dependence and alleviation of the hyperparameter sensitivity problem highlighted by Wade et al., (2023).
2 Logistic-beta processes
2.1 Univariate logistic-beta and Pólya mixing distributions
We begin by reviewing the univariate logistic-beta distribution, which is also referred to as the type IV generalized logistic distribution (Johnson et al.,, 1995) or Fisher’s z distribution (Barndorff-Nielsen et al.,, 1982) in the literature up to location-scale transformations. We say $\eta$ follows a univariate logistic-beta distribution with shape parameters $a,b>0$, if it has density
${\pi_{\textsc{lb}}}(\eta;a,b)=\frac{1}{B(a,b)}\Big{(}\frac{1}{1+e^{-\eta}}\Big{)}^{a}\Big{(}\frac{e^{-\eta}}{1+e^{-\eta}}\Big{)}^{b},\quad\eta\in\mathds{R},$ (1)
where $B(a,b)$ is a beta function. When $a=b=1$, equation (1) reduces to the standard logistic distribution. Applying a logistic transformation $\sigma(x)=1/(1+e^{-x})$ to a logistic-beta random variable yields a beta distribution $\sigma(\eta)\sim\mathrm{Beta}(a,b)$, hence the name logistic-beta. We refer to Johnson et al., (1995, §23.10) for further distributional details.
One important property of the logistic-beta distribution is its representation as a normal variance-mean mixture, as first shown by Barndorff-Nielsen et al., (1982). In particular, the density in equation (1) can be equivalently written as
${\pi_{\textsc{lb}}}(\eta;a,b)=\int_{0}^{\infty}N_{1}\left\{\eta;0.5\lambda(a-b),\lambda\right\}{\pi_{\textsc{po}}}(\lambda;a,b)\mathrm{d}\lambda,$ (2)
where $N_{d}(\mu,\Sigma)$ and $N_{d}(\eta;\mu,\Sigma)$ denote the distribution and density of a $d$-dimensional multivariate normal with mean $\mu$ and covariance $\Sigma$, and $\pi_{\textsc{po}}(\lambda;a,b)$ denotes the density of a Pólya distribution with shape parameters $a$ and $b$. This mixture representation is occasionally employed in regularized logistic regression problems (Gramacy and Polson,, 2012; Polson and Scott,, 2013). Figure 1 shows density plots of the logistic-beta and corresponding Pólya mixing distributions.
Figure 1: Density plot of univariate logistic-beta and associated Pólya mixing distributions.
The Pólya mixing distribution, denoted as $\lambda\sim\mathrm{Polya}(a,b)$, is defined as an infinite sum of exponential distributions $\sum_{k=0}^{\infty}2\epsilon_{k}/\{(k+a)(k+b)\}$, with $\epsilon_{k}$ following standard exponential distributions independently for $k=0,1,\dots$. A truncated version of this expression can be used for approximate random variable generation. The density function of a Pólya can be written as an alternating series (Barndorff-Nielsen et al.,, 1982)
${\pi_{\textsc{po}}}(\lambda;a,b)=\sum_{k=0}^{\infty}(-1)^{k}\binom{a+b+k-1}{k}\frac{k+(a+b)/2}{B(a,b)}\exp\left\{-\frac{(k+a)(k+b)}{2}\lambda\right\}.$ (3)
In practice, this representation brings numerical issues in evaluating the density, especially near the origin where the exponential term in equation (3) has a slow decay. However, we derive a simple identity between Pólya densities among parameter pairs $(a,b)$ that have fixed sum, which we use to develop posterior inference algorithms that avoid direct evaluation of equation (3).
Proposition 1.
The Pólya density in equation (3) satisfies the following identity:
${\pi_{\textsc{po}}}(\lambda;a^{\prime},b^{\prime})=B(a,b)B(a^{\prime},b^{\prime})^{-1}\exp\{\lambda(ab-a^{\prime}b^{\prime})/2\}{\pi_{\textsc{po}}}(\lambda;a,b),$ (4)
where $(a^{\prime},b^{\prime})$ is any pair that satisfies $a^{\prime}+b^{\prime}=a+b$ and $a^{\prime},b^{\prime}>0$.
See Appendix B for all proofs and additional details on the Pólya distribution.
2.2 Multivariate logistic-beta distributions
Based on the normal variance-mean mixture representation in equation (2), we introduce a multivariate logistic-beta distribution that has the same univariate logistic-beta marginals.
Definition 1.
Let $R$ be an $n\times n$ positive semidefinite correlation matrix. We say $\eta$ follows an $n$-dimensional multivariate logistic-beta distribution with shape parameters $a,b>0$ and correlation parameter $R$, denoted as $\eta\sim{\textsc{lb}}(a,b,R)$, if
$\eta\mid\lambda\sim N_{n}\left\{0.5\lambda(a-b)1_{n},\lambda R\right\},\quad\lambda\sim\mathrm{Polya}(a,b),$ (5)
where $1_{n}=(1,\dots,1)^{\mathrm{\scriptscriptstyle T}}\in\mathds{R}^{n}$.
Elementwise logistic transformation of a multivariate logistic-beta distribution induces a multivariate distribution on the unit hypercube $[0,1]^{n}$ with $\mathrm{Beta}(a,b)$ marginals. To see this, since $R$ has a unit diagonal, each component $\eta_{i}$ is marginally logistic-beta distributed with shape parameters $a$ and $b$ for $i=1,\dots,n$. Thus, logistic transformation $\{\sigma(\eta_{1}),\dots,\sigma(\eta_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows a multivariate beta distribution with $R$ controlling dependence. See Fig. 2 for illustrations. Such a multivariate generalization is mentioned in Barndorff-Nielsen et al., (1982), and the density is studied in Grigelionis, (2008). However, there is no literature using this construction for multivariate beta modeling. Alternative multivariate constructions that are related to the logistic-beta are proposed by Bradley et al., (2019) for conjugate spatio-temporal modeling of multinomial data and by Kowal et al., (2019) for temporal dependence in shrinkage.
Figure 2: Density plots of bivariate logistic-beta distributions with different correlation parameters $R_{12}\in\{-0.8,0,0.8\}$ and their logistic transformations. The logistic transformed random variables all have beta marginals with shape parameters $a=2$ and $b=4$.
The main advantage of this construction is that the correlation parameter $R$ explicitly controls dependence with great flexibility. In Proposition 2 and Corollary 1, we present the first two moments of the multivariate logistic-beta and the range of pairwise correlation, illustrating how $R$ controls the strength of linear dependence.
Proposition 2.
Let $\eta\sim{\textsc{lb}}(a,b,R)$ with $R_{ij}$ being the $(i,j)$th element of $R$. Then, we have $E(\eta_{i})=\psi(a)-\psi(b)$, $\operatorname*{var}(\eta_{i})=\psi^{\prime}(a)+\psi^{\prime}(b)$ for $i=1,\dots,n$, and
$\operatorname*{cov}(\eta_{i},\eta_{j})=\begin{cases}2\psi^{\prime}(a)R_{ij},&\text{if }a=b,\\ \psi^{\prime}(a)+\psi^{\prime}(b)+2(R_{ij}-1)\{\psi(a)-\psi(b)\}/(a-b),&\text{if }a\neq b,\end{cases}$
where $\psi(x)$ and $\psi^{\prime}(x)$ denote the digamma and the trigamma functions, respectively.
Corollary 1 (Correlation range).
Let $\eta\sim{\textsc{lb}}(a,b,R)$ with some $2\times 2$ positive semidefinite correlation matrix $R$. If $a=b$, $\operatorname*{corr}(\eta_{1},\eta_{2})$ has a full range $[-1,1]$, and if $a\neq b$, the range of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is $[1-4\{\psi(a)-\psi(b)\}/[(a-b)\{\psi^{\prime}(a)+\psi^{\prime}(b)\}],1]$.
From Proposition 2, the covariance between $\eta_{i}$ and $\eta_{j}$ is a linear function of $R_{ij}$. The Pólya mixing variable $\lambda$ is shared across dimensions so that $R_{ij}=0$ does not imply independence between $\eta_{i}$ and $\eta_{j}$. The covariance between the logistic transformed variables $\operatorname*{cov}\{\sigma(\eta_{i}),\sigma(\eta_{j})\}$ lacks an analytical form, but Monte Carlo estimates are easily obtained. While Corollary 1 shows that arbitrarily high positive correlation can be captured for any choice of $(a,b)$, we have a nontrivial lower bound when $a\neq b$. The restriction on the strength of negative correlation is due to the shared asymmetric marginals (Joe,, 2006).
2.3 Logistic-beta processes and correlation kernels
We define a logistic-beta process on a discrete or continuous domain $\mathscr{X}$.
Definition 2.
Let $\mathcal{R}:\mathscr{X}\times\mathscr{X}\to[-1,1]$ be a correlation kernel which is positive semidefinite with $\mathcal{R}(x,x)=1$ for any $x\in\mathscr{X}$. We say $\{\eta(x)\in\mathds{R}:x\in\mathscr{X}\}$ is a logistic-beta process with shape parameters $a,b>0$ and correlation kernel $\mathcal{R}$, denoted as $\eta(\cdot)\sim{\textsc{lbp}}(a,b,\mathcal{R})$, if every finite collection $\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ follows an $n$-dimensional multivariate logistic-beta with shape parameters $a,b$ and correlation parameter $R$ with $(i,j)$th element $R_{ij}=\mathcal{R}(x_{i},x_{j})$.
From the definition of multivariate logistic-beta (5) that shares a common Pólya mixing variable and properties of the multivariate normal, the logistic-beta process is closed under marginalization and thus is a valid stochastic process as for other mixtures of Gaussian processes (Yu et al.,, 2007). Specifically when $a=b$, it belongs to the family of elliptical processes (Shah et al.,, 2014). The logistic transformation $\sigma\{\eta(\cdot)\}$ has marginals distributed as $\mathrm{Beta}(a,b)$, with dependence induced by the correlation kernel $\mathcal{R}$. Below we provide examples of correlation kernel constructions.
The logistic-beta process can capture groupwise, temporal, and spatial dependence in discrete or continuous domains. When $\mathscr{X}$ corresponds to discrete time indices, we can consider an autoregressive order one ar(1) correlation kernel $\mathcal{R}(x,x^{\prime})=\rho^{|x-x^{\prime}|}$ for $|\rho|<1$. When $\mathscr{X}$ is a continuous spatial domain, a prominent example is the Matérn kernel $\mathcal{R}_{M}(x,x^{\prime};\varrho,\nu)=2^{1-\nu}\Gamma(\nu)^{-1}(\|x-x^{\prime}\|/\varrho)^{\nu}K_{\nu}(\|x-x^{\prime}\|/\varrho)$ with parameters $\varrho,\nu>0$, where $K_{\nu}$ is the modified Bessel function of the second kind; see Fig. 3 for realizations after logistic transformation.
We can consider a more general construction of a correlation kernel $\mathcal{R}$. Let $\phi:\mathscr{X}\to\mathds{R}^{q}$ be a normalized feature map $\phi(x)=\{\phi_{1}(x),\dots,\phi_{q}(x)\}^{\mathrm{\scriptscriptstyle T}}$ such that $\|\phi(x)\|_{2}^{2}=1$ (Graf et al.,, 2003). We assume finite $q$, although infinite extensions are possible. Such a normalized feature map can be obtained from an unnormalized feature map $\varphi(x)$ with a bounded norm by letting $\phi(x)=\varphi(x)/\|\varphi(x)\|_{2}$; for example, $\varphi(x)$ can be spline basis functions for bounded domain $\mathscr{X}$, which induces nonlinear dependence in $x$. Then, a valid correlation kernel $\mathcal{R}$ can be defined as an inner product $\mathcal{R}(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle.$ Such kernels lead to convenient representations of the logistic-beta process presented in Proposition 3, which are helpful in understanding the induced dependence structure and facilitating posterior inference.
Proposition 3.
Consider a normalized feature map $\phi:\mathscr{X}\to\mathds{R}^{q}$ and the corresponding kernel $\mathcal{R}(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle$. Let $\Phi=\{\phi_{k}(x_{i})\}_{ik}$ be an $n\times q$ basis matrix with $i$th row $\phi(x_{i})$. Then, a realization $\eta=\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ from ${\textsc{lbp}}(a,b,\mathcal{R})$ can be equivalently written as the linear predictor representation (6) and the hierarchical representation (7),
$\eta$ =$=\{\psi(a)-\psi(b)\}(1_{n}-\Phi 1_{q})+\Phi\beta,\quad\beta\sim{\textsc{lb}}(a,b,I_{q}),$ (6)
$\eta$ $=0.5\lambda(a-b)1_{n}+\lambda^{1/2}\Phi\gamma,\quad\lambda\sim\mathrm{Polya}(a,b),\quad\gamma\sim N_{q}(0_{q},I_{q}).$ (7)
Representation (6) resembles basis function representations of a Gaussian process with normal priors on coefficients (Gelman et al.,, 2015, §21.1), but with a $q$-dimensional multivariate logistic-beta prior on the coefficients. When we consider an independent Bernoulli response model with logit link $z(x)\sim\mathrm{Ber}[\sigma\{\eta(x)\}]$, equation (6) corresponds to the linear predictor of the logistic regression model with an additional fixed varying intercept term. The hierarchical model representation (7) is useful for posterior inference. Conditional on $\lambda$, the logistic-beta process with the normalized feature map kernel is parameterized with normal coefficients $\gamma$, enabling conditionally conjugate updates.
3 Latent logistic-beta process model for nonparametric binary regression
3.1 Model
We consider a nonparametric binary regression problem in which logit-transformed probabilities of binary responses have a logistic-beta process prior. We highlight advantages of this prior in simplifying posterior inference and prediction, while facilitating extensions. Let $z(x)\in\{0,1\}$ be binary data indexed by $x\in\mathscr{X}$. To induce a $\mathrm{Beta}(a,b)$ prior for $\operatorname*{pr}\{z(x)=1\}$ for each $x$, while enabling inference on probabilities at arbitrary $x^{*}$ values, we choose
$z(x_{i})\mid\eta(x_{i})$ $\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Ber}[\sigma\{\eta(x_{i})\}]\quad(i=1,\dots,n),\quad\eta(\cdot)\sim{\textsc{lbp}}(a,b,\mathcal{R}).$ (8)
By logistic-beta properties, the prior probabilities are marginally $\mathrm{Beta}(a,b)$ distributed, providing a natural extension from the independent beta-Bernoulli model to a dependent process model. The right panel of Fig. 3 shows an example of posterior prediction. We simulated 600 binary data points with $\operatorname*{pr}\{z(x)=1\}=\sigma\{\cos(\pi x)\}$, and fitted the latent logistic-beta process model (8) with $a=2,b=4$ and Matérn correlation kernel with fixed parameters $(\varrho,\nu)=(0.3,1.5)$. The middle part is slightly biased towards zero and has a wider credible interval, reflecting shrinkage towards the prior and uncertainty due to sparse data.
Figure 3: Illustration of the logistic-beta process on $\mathscr{X}=[0,3]$ with Matérn correlation kernel. (Left) Eight realizations of $\sigma\{\eta(x)\}$ with $\eta(\cdot)\sim{\textsc{lbp}}(2,4,\mathcal{R}_{M})$, following $\mathrm{Beta}(2,4)$ marginally. (Right) Posterior mean (solid) and 95% credible interval of latent success probabilities given the binary data $z(x_{i})\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathrm{Ber}[\sigma\{\eta(x_{i})\}]$, $i=1,\dots,n$ represented as small vertical bars. The dashed line corresponds to the true success probabilities.
3.2 Posterior computation
We describe a Markov chain Monte Carlo algorithm for model (8). The key components are Pólya-Gamma data augmentation (Polson et al.,, 2013) and the normal variance-mean mixture representation (5), which leads to a latent Gaussian model conditional on the Pólya mixing variable $\lambda$. In Algorithm 9, we describe a blocked Gibbs sampler. We use notations $z_{1:n}=\{z(x_{1}),\dots,z(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$, $\omega=(\omega_{1},\dots,\omega_{n})^{\mathrm{\scriptscriptstyle T}}$ and $\Omega=\hbox{diag}(\omega)$ for Pólya-Gamma variables, $\eta=(\eta_{1},\dots,\eta_{n})^{\mathrm{\scriptscriptstyle T}}=\{\eta(x_{1}),\dots,\eta(x_{n})\}^{\mathrm{\scriptscriptstyle T}}$ and $R=\{\mathcal{R}(x_{i},x_{j})\}_{i,j}$. Here we assume that the correlation kernel $\mathcal{R}$ is known for conciseness, but incorporating unknown correlation kernel parameters is straightforward via an additional sampling step.
[1] Sample $\omega_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\pi(\omega_{i}\mid\eta_{i})$ for $i=1,\dots,n$, which is Pólya-Gamma with parameters $1$ and $\eta_{i}$.
[2] Sample $\lambda$ from $\pi(\lambda\mid\omega,z_{1:n})$, where $\eta$ is marginalized out,
$\pi(\lambda\mid\omega,z_{1:n})\propto{\pi_{\textsc{po}}}(\lambda;a,b)\times N_{n}\big{\{}\Omega^{-1}(z_{1:n}-0.51_{n});0.5\lambda(a-b)1_{n},\lambda R+\Omega^{-1}\big{\}}.$ (9)
[3] Sample $\eta$ from $\pi(\eta\mid\lambda,\omega,z_{1:n})\propto\pi(z_{1:n}\mid\eta,\omega)\times\pi(\eta\mid\lambda)$, which is multivariate normal.
Algorithm 1 One cycle of a blocked Gibbs sampler for model (8).
Step 1 of Algorithm 9 updates Pólya-Gamma auxiliary variables $\omega$ using a rejection sampler (Polson et al.,, 2013). Data augmentation with Pólya-Gamma auxiliary variables converts the Bernoulli likelihood (8) into a normal likelihood in terms of logit probabilities $\eta$ conditional on Pólya-Gamma variables $\omega$. Leveraging (5), $\pi(\eta\mid\lambda)$ is also a multivariate normal in terms of $\eta$. This leads to joint conditional $\pi(\eta,\lambda\mid\omega,z_{1:n})\propto\pi(z_{1:n}\mid\eta,\omega)\pi(\eta\mid\lambda)\pi(\lambda)$ that is multivariate normal in terms of $\eta$. Thus, the computation of collapsed conditional $\pi(\lambda\mid\omega,z_{1:n})=\int\pi(\lambda,\eta\mid\omega,z_{1:n})\mathrm{d}\eta$ can integrate out $\eta$ by exploiting normal-normal conjugacy, greatly improving mixing behavior. Furthermore, again by conjugacy, Step 3 of Algorithm 9 updates $\eta$ from a multivariate normal; for example when correlation matrix $R$ is full rank,
$\eta\mid\omega,\lambda,z_{1:n}\sim N_{n}\big{[}(\Omega+\lambda^{-1}R^{-1})^{-1}\{(z_{1:n}-0.51_{n})+0.5(a-b)R^{-1}1_{n}\},(\Omega+\lambda^{-1}R^{-1})^{-1}\big{]}.$
Thus, the logistic-beta process offers significant computational benefits compared to multivariate beta constructions that rely on coordinate-wise updating (Trippa et al.,, 2011; DeYoreo and Kottas,, 2018) or Metropolis-Hastings in high-dimensional space (Arbel et al.,, 2016). Sampling $\lambda$ from a collapsed conditional density in Step 2 remains nontrivial and involves the Pólya density ${\pi_{\textsc{po}}}$ that is numerically unstable to evaluate; in Appendix A we propose a novel class of adaptive Pólya proposals that bypass evaluation of the Pólya density based on Proposition 1. In Appendix C, we present further computational strategies based on scalable Gaussian process methods that preserve marginal variances, including normalized feature map kernel with $q\ll n$; the representations (6) and (7) both highlight the dimension reduction from $n$ to $q$.
4 Logistic-beta dependent Dirichlet processes
4.1 Definition and prior properties
The latent logistic-beta process model can be easily employed in more complex Bayesian nonparametric models. For example, we introduce a logistic-beta dependent Dirichlet process for modeling dependent random probability measures. Random probability measures $\{G_{x}:x\in\mathscr{X}\}$ follow a logistic-beta dependent Dirichlet process with concentration $b$, correlation kernel $\mathcal{R}$, and independent atom processes $\{\theta_{h}(x):x\in\mathscr{X}\}$, $h\in\mathds{N}$ with $G^{0}_{x}$ the base measure, if
$G_{x}(\cdot)=\sum_{h=1}^{\infty}\Big{(}\sigma\{\eta_{h}(x)\}\prod_{l<h}\left[1-\sigma\{\eta_{l}(x)\}\right]\Big{)}\delta_{\theta_{h}(x)}(\cdot),\,\,\,\,\eta_{h}(\cdot)\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{\textsc{lbp}}(1,b,\mathcal{R})\,\,(h\in\mathds{N}).$ (10)
Since the stick-breaking ratios $\sigma\{\eta_{h}(x)\}$ have $\mathrm{Beta}(1,b)$ marginals for any $x\in\mathscr{X}$ and are independent across $h=1,2\dots$, the random probability measure $G_{x}$ for any given $x\in\mathscr{X}$ marginally follows a Dirichlet process with concentration parameter $b$. The hyperparameter $b$ has a familiar interpretation, aiding in elicitation.
To gain a deeper understanding of the dependence structure induced by the logistic-beta dependent Dirichlet process, we study the correlation between random probability measures, focusing on the dependence induced by the weights by assuming an atom process that does not depend on $x$, referred to as a single-atoms process.
Theorem 1.
Consider a single-atoms logistic-beta dependent Dirichlet process $\{G_{x}:x\in\mathscr{X}\}$ with concentration parameter $b$ and correlation kernel $\mathcal{R}$, where its atoms are identically and independently distributed from a nonatomic base measure. Let $\mu(x_{i},x_{j})=E[\sigma\{\eta(x_{i})\}\sigma\{\eta(x_{j})\}]$ with $\eta(\cdot)\sim{\textsc{lbp}}(1,b,\mathcal{R})$. Then, for any Borel set $B$,
$\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}={(1+b)^{2}}/\left\{2\mu(x_{i},x_{j})^{-1}-(1+b)\right\}.$ (11)
The correlation (11) does not depend on the choice of $B$ and attains a maximum of $1$ when $x_{i}=x_{j}$. The minimum correlation is achieved when $\mu(x_{i},x_{j})$ is minimized, whose value depends on $b$ due to the nontrivial correlation lower bound described in Corollary 1.
To investigate the range of dependence, we analyze and compare the greatest lower bounds of correlation (11) under four different single-atoms dependent Dirichlet process models based on dependent beta stick-breaking ratios: (M1) logistic-beta dependent Dirichlet process; (M2) The process proposed by DeYoreo and Kottas, (2018); (M3) The process proposed by Nieto-Barajas et al., (2012); (M4) Copula-based dependent Dirichlet process. For M4, the minimum correlation is achieved with counter-monotonic beta random variables; examples include Gaussian copula with perfect negative correlation. Under the single-atoms processes, where atoms are shared across $x$, the correlation (11) cannot be negative even though the correlation between weights is negative (Ascolani et al.,, 2023). For completeness, we add brief descriptions of M2–M4 and their correlation lower bounds in Appendix B.3.
Figure 4: Range of dependence induced by single-atoms dependent Dirichlet processes (M1)–(M4) against the concentration parameter $b$ on a $\log_{2}$ scale. (Left) Infimum of $\operatorname*{corr}\{V_{h}(x_{i}),V_{h}(x_{j})\}$, the correlation between $\mathrm{Beta}(1,b)$ stick-breaking variables. (Right) Infimum of $\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}$, the correlation between random probability measures evaluated on the same set. All processes (M1)–(M4) attain perfect correlation of 1 when $x_{i}=x_{j}$.
Figure 4 illustrates the range of dependence that processes M1–M4 can capture, with M4 being the widest. The behaviour of the minimum correlation of M1 varies distinctly under different values of $b$ compared to others. For small to moderate $b$, M1 can capture a much wider range of dependence compared to M2 and M3 by allowing negatively correlated stick-breaking ratios. When $b$ is large, the correlation range of M1 becomes limited due to the nontrivial lower bound of correlation described in Corollary 1, and it gives useful guidance on the choice of $b$. In practice, we suggest choosing $b$ less than 8 for M1 so that a negative correlation between beta stick-breaking variables is possible.
4.2 Logistic-beta dependent Dirichlet process mixture model
We model a covariate-dependent response density as $f(y\mid x)=\int\mathcal{K}(y\mid\theta)\mathrm{d}G_{x}(\theta)$, where $\{G_{x}\}$ follows a logistic-beta dependent Dirichlet process and $\mathcal{K}(\cdot\mid\theta)$ is a mixture component distribution parametrized by $\theta$. For example, we choose a normal kernel $\mathcal{K}(y\mid\theta)=N_{1}(y;\mu,\tau^{-1})$ and atom processes $\theta_{h}(x)=\{\mu_{h}(x),\tau_{h}(x)\}$ with $\mu_{h}(x)=\beta_{0h}+\beta_{1h}x$, $\tau_{h}(x)=\tau_{h}$,
$f(y_{i}\mid x_{i})$ $=\sum_{h=1}^{\infty}\Big{(}\sigma\{\eta_{h}(x_{i})\}\prod_{l<h}[1-\sigma\{\eta_{l}(x_{i})\}]\Big{)}N_{1}(y_{i};\mu_{h}(x_{i}),\tau_{h}^{-1})\quad(i=1,\dots,n),$ (12)
$\eta_{h}(\cdot)$ $\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}{\textsc{lbp}}(1,b,\mathcal{R}),\,\,\,\,(\beta_{0h},\beta_{1h})^{\mathrm{\scriptscriptstyle T}}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N_{2}(0_{2},\Sigma_{\beta}),\quad\tau_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Ga}(a_{\tau},b_{\tau})\,\,\,\,(h\in\mathds{N}),$ (13)
where logistic-beta processes in (13) and atom processes both induce dependence on $x$, leading to a mixture of normal linear regression models with covariate-dependent weights.
To adapt Algorithm 9 to model (12)-(13), we rely on the continuation-ratio logit representation of the multinomial distribution (Tutz,, 1991), which is also used in other stick-breaking process mixture models (Rodríguez and Dunson,, 2011; Rigon and Durante,, 2021). Denoting the mixture component allocations as $\{s_{i}\}_{i=1}^{n}$ and $\mathscr{Z}_{h}=\{z_{ih}:s_{i}>h-1\}$ as a set of binary variables such that $z_{ih}=1$ if $s_{i}=h$ and $z_{ih}=0$ if $s_{i}>h$, $\mathscr{Z}_{h}$ corresponds to observations of the latent logistic-beta process model for binary data with indices $I_{h}=\{i:s_{i}>h-1\}$, independently across $h=1,2,\dots$. We work on a truncated version having maximum number of components $H$ with $\sigma\{\eta_{H}(\cdot)\}\equiv 1$; if all the components are occupied we recommend increasing $H$.
[1] Sample component allocations $s_{i}\in\{1,\dots,H\}$ for $i=1,\dots,n$,
$\operatorname*{pr}(s_{i}=h\mid-)\propto\sigma\{\eta_{h}(x_{i})\}\prod_{l=1}^{h-1}[1-\sigma\{\eta_{l}(x_{i})\}]\mathcal{K}(y_{i}\mid\theta_{h}(x_{i})).$
[2] For $h=1$ to $h=H-1$, with index set $I_{h}=\{i:s_{i}>h-1\}$, $\lambda_{h}$ and $\eta_{h}(\cdot)$,
[0] Run Algorithm 9 with binary data $\mathscr{Z}_{h}=\{\mathds{1}(s_{i}=h):i\in I_{h}$}, update $\lambda_{h}$ and $\eta_{h}(\cdot)$.
[3] Sample component-specific parameters $(\beta_{0h},\beta_{1h},\tau_{h})^{\mathrm{\scriptscriptstyle T}}$ for $h=1,\dots,H$,
$\pi(\beta_{0h},\beta_{1h},\tau_{h}\mid-)\propto G^{0}(\beta_{0h},\beta_{1h},\tau_{h})\prod_{i:s_{i}=h}\mathcal{K}(y_{i}\mid\theta_{h}(x_{i})).$
Algorithm 2 One cycle of the blocked Gibbs sampler for the mixture model (12)–(13).
5 Empirical demonstrations
5.1 Simulation studies
Our simulation studies have two main objectives. Firstly, we aim to assess the benefits of posterior computation strategies involved in latent logistic-beta process models. We focus on how the blocking strategy in the Gibbs sampler and adaptive Pólya proposal affect mixing. Secondly, we compare the latent logistic-beta process with Gaussian copula models, which also give dependent random probabilities with beta marginals. We consider a nonparametric binary regression problem on a spatial domain $\mathscr{X}=[0,1]^{2}$, where spatial coordinates $\{x_{i}\}_{i=1}^{n}$ are generated uniformly at random from $\mathscr{X}$.
We considered the latent logistic-beta process model (8) with $(a,b)=(1,2)$ and the Gaussian copula model as two different data generation models. The Gaussian copula model first draws a sample $\big{\{}\zeta(x_{1}),\dots,\zeta(x_{n})\big{\}}^{\mathrm{\scriptscriptstyle T}}$ from a mean zero unit variance Gaussian process and then applies transformation $\zeta\mapsto F_{B}^{-1}\{F_{Z}(\zeta)\}$ to yield beta marginals, where $F_{Z}$ is the standard normal c.d.f. and $F_{B}^{-1}$ is the inverse c.d.f. of the Beta($1,2$) distribution. For both the latent logistic-beta process and Gaussian copula models, we used Matérn correlation kernels $\mathcal{R}_{M}(\cdot,\cdot;\varrho,\nu=1.5)$. We considered different degrees of spatial dependence $\varrho\in\{0.1,0.2,0.4\}$, resulting in six different data generation scenarios. Given the success probabilities, binary data $\{z(x_{i})\}_{i=1}^{n}$ of size $n=500$ were independently sampled, with $n_{\mathrm{test}}=100$ data points reserved to evaluate predictive performance. We simulated 100 replicated datasets for each data generation scenario.
The latent logistic-beta process and Gaussian copula models are fitted to binary data with size $n_{\mathrm{train}}=400$. Beta parameters are fixed at the true value $(a,b)=(1,2)$, and Matérn correlation kernels are adopted with a fixed smoothness parameter $\nu=1.5$. The Gibbs sampler for the latent logistic-beta process model is implemented in R (R Core Team,, 2023), and we explore four posterior inference algorithms, considering blocking and/or the adaptive Pólya proposal scheme in Step 2 of Algorithm 9. For the Gaussian copula model, we use Stan (Carpenter et al.,, 2017) to implement Hamiltonian Monte Carlo. For both models, we run 2,000 Markov chain Monte Carlo iterations and record wall-clock running time, with the first 1,000 samples discarded as a burn-in. Posterior predictive distributions of success probabilities in the test dataset are compared with true values to assess predictive performance.
Table 1: Simulation results under various posterior inference algorithm settings. Entries are averages over 100 replicates, with Monte Carlo standard errors shown in parentheses. LBP, logistic-beta process; ESS, effective sample size of $\lambda$ based on 1000 posterior samples; ESS/sec, ESS divided by running time in seconds.
Data generation Algorithm settings ESS ESS/sec % Acceptance rate
Latent LBP, $\varrho=0.1$ Blocked Adapted 245.08 (12.86) 3.35 (0.18) 54.28 (1.03)
Non-adapted 196.35 (11.28) 2.69 (0.15) 49.39 (1.39)
Non-blocked Adapted 7.89 (0.35) 0.14 (0.01) 13.60 (0.26)
Non-adapted 7.13 (0.37) 0.13 (0.01) 12.44 (0.31)
Latent LBP, $\varrho=0.2$ Blocked Adapted 257.01 (16.32) 2.89 (0.18) 62.26 (1.12)
Non-Adapted 247.83 (16.53) 2.82 (0.19) 57.60 (1.47)
Non-blocked Adapted 7.31 (0.32) 0.11 (0.00) 13.42 (0.22)
Non-adapted 6.58 (0.32) 0.10 (0.00) 12.48 (0.31)
Latent LBP, $\varrho=0.4$ Blocked Adapted 368.26 (17.54) 4.12 (0.20) 66.12 (0.98)
Non-adapted 328.72 (19.00) 3.67 (0.21) 61.45 (1.32)
Non-blocked Adapted 6.40 (0.29) 0.09 (0.00) 13.01 (0.22)
Non-adapted 6.56 (0.32) 0.10 (0.00) 12.88 (0.31)
Table 1 reports the performance of four different latent logistic-beta process posterior inference algorithm configurations based on the mixing behaviour of $\lambda$. “Non-blocked” refers to sampling $\lambda$ from $\pi(\lambda\mid\omega,\eta,z_{1:n})$ instead of $\pi(\lambda\mid\omega,z_{1:n})$, and “non-adapted” corresponds to the fixed independent Metropolis-Hastings proposal distribution of $\lambda$ using prior distribution $\mathrm{Polya}(a,b)$. We consider effective sample size based on 1,000 posterior samples, effective sample rate, and an acceptance rate of $\lambda$ from an independent Metropolis-Hastings sampler where a higher acceptance rate is desirable (Robert and Casella,, 2004, §7.6.1). The blocking scheme has a substantial impact on the mixing behaviour, leading to an effective sample size increase of up to 50 times. The adaptive Pólya proposal scheme improves the acceptance rate of $\lambda$ in all scenarios.
Table 2: Simulation results comparing latent logistic-beta process and Gaussian copula models. Entries are averaged over 100 replicates, with Monte Carlo standard errors in parentheses. LBP, logistic-beta process; RMSE, root mean squared error; CRPS, continuous ranked probability score; mESS/sec, multivariate effective sample size of latent success probabilities divided by running time in seconds.
Data generation Model RMSE ×$\times 100$ mean CRPS ×$\times 100$ mESS/sec
training test training test
Gauss. copula, $\varrho=0.1$ Latent LBP 11.93 (0.14) 12.32 (0.17) 6.59 (0.10) 6.80 (0.11) 21.11 (0.21)
Gauss. copula 11.82 (0.13) 12.24 (0.16) 6.48 (0.09) 6.71 (0.11) 0.48 (0.01)
Gauss. copula, $\varrho=0.2$ Latent LBP 8.67 (0.15) 8.80 (0.16) 4.78 (0.10) 4.85 (0.10) 17.58 (0.16)
Gauss. copula 8.61 (0.16) 8.75 (0.17) 4.74 (0.10) 4.82 (0.11) 0.40 (0.01)
Gauss. copula, $\varrho=0.4$ Latent LBP 6.11 (0.16) 6.14 (0.16) 3.39 (0.10) 3.41 (0.10) 17.41 (0.17)
Gauss. copula 6.10 (0.16) 6.13 (0.16) 3.38 (0.10) 3.40 (0.10) 0.47 (0.01)
Table 2 reports the in-sample and predictive performance of the latent logistic-beta process and Gaussian copula models, as well as the sampling efficiency with Gaussian copula data generation scenarios. In terms of root mean square error and mean continuous ranked probability score of latent success probabilities across both training and test data, the models produce nearly identical results. Although data were generated from the Gaussian copula model, the latent logistic-beta process model is able to recover latent probabilities as accurately as the true model.
Regarding sampling efficiency of latent success probabilities measured by multivariate effective sample size (Vats et al.,, 2019) divided by the algorithm’s running time in seconds, the latent logistic-beta process model dramatically outperforms the Gaussian copula model due to the significantly longer running time of the Gaussian copula model. Although the implementation of the Gaussian copula model with the Stan program is based on C++ code with highly optimized algorithms, the computation of the log-posterior and its gradient at each step demands high computational costs. The algorithm of the latent logistic-beta process model is currently implemented using R software, which has the potential for improvement in running time with a lower-level programming language. Appendix D contains further details on simulation settings and additional simulation results under different data generation scenarios.
5.2 Application to pregnancy outcome data analysis
We use the logistic-beta dependent Dirichlet process mixture model (12)–(13) to analyze DDE exposure and pregnancy outcome data (Longnecker et al.,, 2001). DDE is a metabolite of chemical compound DDT that was widely used as an insecticide. We analyze how the conditional distribution of gestational age of delivery $y$ changes with DDE exposure level $x$, focusing on the probability of preterm birth, defined as gestational age of delivery less than 37 weeks. The dataset is publicly available in R package BNPmix (Corradin et al.,, 2021). For illustration we focus on the $n=1023$ women who smoked during pregnancy; results for other groups are in Appendix D.2.
Since there is no publicly available software for a dependent Dirichlet process with covariate-dependent weights, we compare with the logit stick-breaking process (Ren et al.,, 2011). We choose basis functions as in Rigon and Durante, (2021) by considering normalized feature map kernels obtained from a natural cubic spline basis with 6 degrees of freedom. In both our model and logit stick-breaking, we use a mixture of linear regressions with covariate-dependent weights as in (12). We compare the results in terms of the probability of preterm birth given DDE exposure. For the logistic-beta dependent Dirichlet process, we consider three different settings of parameter $b\in\{1/5,1,2\}$, referred to as settings 1, 2, and 3, respectively. This corresponds to prior co-clustering probabilities for any $x$ of $5/6$, $1/2$, and $1/3$. For the logit stick-breaking process, we choose hyperparameters such that prior co-clustering probabilities closely match with the proposed model within each setting.
Posterior computation is carried out with Gibbs samplers for both models. We obtained 35,000 iterations, with the first 5,000 samples discarded as burn-in. Both algorithms were implemented in R and had similar running times; 7.5 minutes for the logistic-beta dependent Dirichlet process model and 5.6 minutes for the logit stick-breaking process model under setting 2.
Figure 5: Estimated probability of preterm birth with 95% credible intervals against DDE exposure level from the logistic-beta dependent Dirichlet process model (top) and the logit stick-breaking process model (bottom) with three different hyperparameter settings. DDE exposure data are represented by small vertical bars.
The estimated preterm birth probabilities given DDE exposures are summarized in Fig. 5. For both models, the probabilities tend to increase with DDE levels, with wider credible intervals at high levels due to the increasingly sparse data. There is a stark difference between the two models in terms of how the estimates and credible intervals vary according to hyperparameter choice. Although each hyperparameter setting yields similar prior properties and both models use the same basis functions to incorporate dependence in $x$, Fig. 5 shows that the posterior estimate of the logit stick-breaking process is highly sensitive compared to the proposed model. Specifically, the hyperparameter choice of the logit stick-breaking process model has a substantial impact on uncertainty estimates at high DDE exposure levels, where setting 1 and setting 3 of the logit stick-breaking process model have much wider and narrower credible intervals, respectively, compared to the proposed model. This sensitivity, as noted in Wade et al., (2023), makes conclusions drawn from the logit stick-breaking process model questionable, whereas the proposed logistic-beta dependent Dirichlet process model is much more robust to the choice of hyperparameters.
Appendix D.2 contains details on pregnancy outcome data analysis settings and additional results with the non-smoking group of size $n=1290$ that gives the same conclusion.
6 Discussion
Although we have focused on two particular applications, logistic-beta processes can be usefully integrated within many other Bayesian nonparametric models. A particularly interesting case is to incorporate covariate dependence into Indian buffet processes, relying on a beta-Bernoulli representation. An interesting application is to species distribution modeling in ecology (Elith and Leathwick,, 2009), where the features correspond to species and covariates correspond to environmental data, such as location or temperature. Such a model can accommodate and predict new species discovery, which is common in ecological studies of insect and fungi.
Acknowledgements
The research of Changwoo Lee and Huiyan Sang was partially supported by the U.S. National Science Foundation (2210456 and 2220231). The research of Alessandro Zito and David Dunson was supported by U.S. Office of Naval Research (N00014-21-1-2510-P00004-7), and National Institutes of Health Grant (R01ES035625). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (856506).
Supplementary Material
Codes for reproducing all analyses are available at the following Github repository: https://github.com/changwoo-lee/logisticbeta-reproduce.
Appendices
Appendix A Details of adaptive Pólya proposal
We describe the adaptive Pólya proposal scheme for sampling $\lambda$ from a distribution that involves Pólya prior density ${\pi_{\textsc{po}}}(\lambda;a,b)$, such as expression (9) in Step 2 of Algorithm 9. The evaluation of Pólya density ${\pi_{\textsc{po}}}$ is problematic since it is expressed in terms of alternating series whose evaluation is numerically unstable; see equation (3). In the context of regularized logistic regression where the same problem arises, Gramacy and Polson, (2012) proposed using independent Metropolis-Hastings with Pólya prior proposal $\mathrm{Polya}(a,b)$ to avoid the evaluation of ${\pi_{\textsc{po}}}$ as these cancel out in the acceptance ratio. However, such a sampler ignores all information in the data and can suffer from a low acceptance rate.
The adaptive Pólya proposal scheme selects the proposal distribution as $\mathrm{Polya}(a^{\prime},b^{\prime})$ with suitably chosen $(a^{\prime},b^{\prime})$ pair such that $a^{\prime}+b^{\prime}=a+b=c$, leveraging the identity in Proposition 1 to avoid evaluation of the Pólya density ${\pi_{\textsc{po}}}$. To see this, for example with Metropolis-Hastings, a new candidate is drawn from $\lambda^{\star}\sim\mathrm{Polya}(a^{\prime},b^{\prime})$, and accepted with probability $\min\{1,\exp\{(\lambda-\lambda^{\star})(ab-a^{\prime}b^{\prime})/2\}g(\lambda^{\star})/g(\lambda)\}$ where ${\pi_{\textsc{po}}}$ terms are cancelled out; here $g(\lambda)$ corresponds to multivariate normal density in (9) as a function of $\lambda$. This adaptive Pólya proposal scheme can be used in more complex transition kernels, such as a particle Gibbs sampler (Andrieu et al.,, 2010), see Appendix C.2 for examples.
There are many ways to incorporate previous samples’ information in the choice of $(a^{\prime},b^{\prime})$. An example is the moment matching method with a running average of $\lambda$, denoted as $\bar{\lambda}$, that is recursively updated as $\bar{\lambda}\leftarrow\bar{\lambda}(m-1)/m+\lambda/m$ in $m$th iteration of the algorithm, leading to adaptive Metropolis-within-Gibbs (Roberts and Rosenthal,, 2009). Assuming $a^{\prime}\leq b^{\prime}$ without loss of generality due to the symmetry of the Pólya distribution in its arguments, which implies $a^{\prime}\leq c/2$, the $\mathrm{Polya}(a^{\prime},c-a^{\prime})$ has a mean $h(a^{\prime})=2\{\psi(a^{\prime})-\psi(c-a^{\prime})\}/(2a^{\prime}-c)$ (or $h(a^{\prime})=2\psi^{\prime}(c/2)$ if $a^{\prime}=c/2$) that is decreasing on $a^{\prime}\in(0,c/2]$. Thus, we set $a^{\prime}=(a+b)/2$ if $\bar{\lambda}$ is less than or equal to $2\psi^{\prime}((a+b)/2)$, and otherwise we set $a^{\prime}$ to the solution of $h(x)=\bar{\lambda}$ which can be easily obtained with numerical optimization algorithms.
Appendix B Proofs and additional propositions
B.1 Properties of Pólya distributions
The Pólya distribution serves as a mixing distribution in the normal variance-mean mixture representation of logistic-beta, which includes standard logistic distribution as a special case. Thus, the Pólya distribution generalizes the squared Kolmogorov–Smirnov distribution (Andrews and Mallows,, 1974), used as the scale mixing distribution in Gaussian scale mixture representations of the standard logistic distribution (Chen and Dey,, 1998; Holmes and Held,, 2006).
We collect results related to the Pólya distribution, including its moments (Proposition B.1) and density identity (Proposition 1).
Proposition B.1.
The mean and variance of $\lambda\sim{\pi_{\textsc{po}}}(a,b)$ are
$E(\lambda)=\begin{cases}2\{\psi(a)-\psi(b)\}/(a-b),&\text{if }a\neq b,\\ 2\psi^{\prime}(a),&\text{if }a=b.\end{cases}$
$\operatorname*{var}(\lambda)=\begin{cases}\frac{4}{(a-b)^{2}}\left[\psi^{\prime}(a)+\psi^{\prime}(b)-\frac{2}{(a-b)}\{\psi(a)-\psi(b)\}\right],&\text{if }a\neq b,\\ 2\psi^{(3)}(a)/3,&\text{if }a=b.\end{cases}$
where $\psi^{(3)}$ is a 3rd derivative of digamma function $\psi$.
Proof.
First check the mean. When $a\neq b$,
$E(\lambda)=\sum_{k=0}^{\infty}\frac{2}{(k+a)(k+b)}$ $=\frac{2}{a-b}\sum_{k=0}^{\infty}\left(\frac{1}{k+b}-\frac{1}{k+a}\right)=\frac{2\{\psi(a)-\psi(b)\}}{a-b},$
see Olver et al., (2010, §5.7.6) using subtraction. When $a=b$, we have $\sum_{k=0}^{\infty}2/(k+a)^{2}=2\psi^{\prime}(a)$, see Olver et al., (2010, §5.15.1). For variance, if $a\neq b$,
$\operatorname*{var}(\lambda)=\sum_{k=0}^{\infty}\frac{4}{(k+a)^{2}(k+b)^{2}}$ $=\frac{4}{(a-b)^{2}}\sum_{k=0}^{\infty}\left\{\frac{1}{(k+a)^{2}}+\frac{1}{(k+b)^{2}}-\frac{2}{(k+a)(k+b)}\right\}$
$=\frac{4}{(a-b)^{2}}\left[\psi^{\prime}(a)+\psi^{\prime}(b)-\frac{2}{(a-b)}\{\psi(a)-\psi(b)\}\right]$
and when $a=b$, $\operatorname*{var}(\lambda)=4(-1)^{3+1}\psi^{(3)}(a)/3!=2\psi^{(3)}(a)/3$; see Olver et al., (2010, §25.11.12). ∎
Next, we present a proof of Proposition 1, an identity between Pólya densities.
Proof.
Let $a,b>0$ be given and $(a^{\prime},b^{\prime})$ be a pair that satisfies $a^{\prime}+b^{\prime}=a+b$ and $a^{\prime},b^{\prime}>0$. Starting from (3),
${\pi_{\textsc{po}}}(\lambda;a^{\prime},b^{\prime})$
$=\sum_{k=0}^{\infty}(-1)^{k}\binom{a^{\prime}+b^{\prime}+k-1}{k}\frac{(k+(a^{\prime}+b^{\prime})/2)}{B(a^{\prime},b^{\prime})}\exp\left(-\frac{(k+a^{\prime})(k+b^{\prime})}{2}\lambda\right)$
$=\frac{B(a,b)}{B(a^{\prime},b^{\prime})}\sum_{k=0}^{\infty}(-1)^{k}\binom{a+b+k-1}{k}\frac{(k+(a+b)/2)}{B(a,b)}\exp\left(-\frac{(k+a)(k+b)}{2}\lambda+\frac{(ab-a^{\prime}b^{\prime})\lambda}{2}\right)$
$=\frac{B(a,b)}{B(a^{\prime},b^{\prime})}\exp\left(\frac{(ab-a^{\prime}b^{\prime})\lambda}{2}\right){\pi_{\textsc{po}}}(\lambda;a,b),$
which proves Proposition 1. ∎
B.2 Properties of multivariate logistic-beta and logistic-beta processes
We summarize results related to multivariate logistic-beta and logistic-beta processes, including the moments (Proposition 2), correlation range (Corollary 1), and representations of logistic-beta processes under normalized feature map kernel (Proposition 3).
Below we present proofs of Proposition 2 and Corollary 1.
Proof.
From Johnson et al., (1995, §23.10), when $\eta$ is a univariate logistic-beta random variable with parameters $a$ and $b$, we have $E(\eta)=\psi(a)-\psi(b)$ and $\operatorname*{var}(\eta)=\psi^{\prime}(a)+\psi^{\prime}(b)$. Let $\eta\sim{\textsc{lb}}(a,b,R)$ with $R_{ij}$ being the $(i,j)$th element of $R$. By the law of total covariance, the covariance and correlation is given by
$\operatorname*{cov}(\eta_{i},\eta_{j})$ $=E\{\operatorname*{cov}(\eta_{i},\eta_{j}\mid\lambda)\}+\operatorname*{cov}\{E(\eta_{i}\mid\lambda),E(\eta_{j}\mid\lambda)\}$
$=R_{ij}E(\lambda)+\operatorname*{cov}\{0.5\lambda(a-b),0.5\lambda(a-b)\}$
=$=\begin{cases}2\psi^{\prime}(a)R_{ij},&a=b,\\ \psi^{\prime}(a)+\psi^{\prime}(b)+2(R_{ij}-1)\left\{\frac{\psi(a)-\psi(b)}{a-b}\right\},&a\neq b.\end{cases}$
and
$\operatorname*{corr}(\eta_{i},\eta_{j})=\begin{cases}R_{ij},&a=b,\\ 1+(R_{ij}-1)\left[\frac{2\{\psi(a)-\psi(b)\}}{(a-b)\{\psi^{\prime}(a)+\psi^{\prime}(b)\}}\right],&a\neq b.\end{cases}$
which shows Proposition 2. Considering a bivariate case, if $a=b$, the range of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is $[-1,1]$. If $a\neq b$, the upper bound 1 of $\operatorname*{corr}(\eta_{1},\eta_{2})$ is attained at $R_{12}=1$, and lower bound of $1-4\{\psi(a)-\psi(b)\}/\{(a-b)(\psi^{\prime}(a)+\psi^{\prime}(b))\}$ is attained at $R_{12}=-1$. This shows Corollary 1. ∎
Next, we present a proof of Proposition 3.
Proof.
Consider a normalized feature map $\phi:\mathscr{X}\to\mathds{R}^{q}$ and the corresponding correlation kernel $\mathcal{R}(x,x^{\prime})=\langle\phi(x),\phi(x^{\prime})\rangle$ for logistic-beta process. Let $\Phi=[\phi_{k}(x_{i})]_{ik}$ be an $n\times q$ basis matrix with $i$th row corresponding to $\phi(x_{i})$. Then, for a realization $\eta=(\eta(x_{1}),\dots,\eta(x_{n}))^{\mathrm{\scriptscriptstyle T}}$, we have
$\eta\mid\lambda\sim N_{n}\left(0.5\lambda(a-b)1_{n},\lambda\Phi\Phi^{\mathrm{\scriptscriptstyle T}}\right),\quad\lambda\sim\mathrm{Polya}(a,b).$
However, $\eta\mid\lambda$ is a rank deficient multivariate normal when $q<n$. By introducing $\gamma\sim N_{q}(0_{q},I_{q})$, it can be equivalently written as
$\eta=0.5\lambda(a-b)1_{n}+\lambda^{1/2}\Phi\gamma,\quad\lambda\sim\mathrm{Polya}(a,b),\quad\gamma\sim N_{q}(0_{q},I_{q}),$
which corresponds to the hierarchical representation of Proposition 3. Also,
$\eta$ $=0.5\lambda(a-b)1_{n}+\lambda^{1/2}\Phi\gamma$
$=0.5\lambda(a-b)1_{n}-0.5\lambda(a-b)\Phi 1_{q}+0.5\lambda(a-b)\Phi 1_{q}+\lambda^{1/2}\Phi\gamma$
$=0.5\lambda(a-b)(1_{n}-\Phi 1_{q})+\Phi\{0.5\lambda(a-b)1_{q}+\lambda^{1/2}\gamma\}.$
Thus, if $\lambda\sim\mathrm{Polya}(a,b)$ is marginalized out, by the linearity of expectation and definition of multivariate logistic-beta, we have
$\eta=\{\psi(a)-\psi(b)\}(1_{n}-\Phi 1_{q})+\Phi\beta,\quad\beta\sim{\textsc{lb}}(a,b,I_{q}),$
where we used $E(\lambda)=2\{\psi(a)-\psi(b)\}/(a-b)$ as in Proposition B.1. This shows the linear predictor representation of Proposition 3. ∎
B.3 Logistic-beta dependent Dirichlet process and related models
We present results on the logistic-beta dependent Dirichlet process, including full weak support property (Theorem B.1) and correlation between random probability measures (Theorem 1). We first present proof of Theorem 1, which analyzes prior correlation structure induced by logistic-beta dependent Dirichlet process.
Proof.
The proof closely follows that for Theorem 1 of Dunson and Park, (2008). Recall that we are considering a single-atoms logistic-beta dependent Dirichlet process, where the atom process does not depend on $x$. Denote the base measure $G_{0}$ and $V_{h}(x)=\sigma\{\eta_{h}(x)\}$, where marginally $V_{h}(x)\sim\mathrm{Beta}(1,b)$. Since the dependent Dirichlet process prior implies $G_{x}$ marginally follows a Dirichlet process with concentration parameter $b$, we have $E\{G_{x}(B)\}=G_{0}(B)$ and $\operatorname*{var}\{G_{x}(B)\}=G_{0}(B)\{1-G_{0}(B)\}/(1+b)$ (Müller et al.,, 2015). In addition to the notation $\mu(x_{i},x_{j})=E[\sigma\{\eta(x_{i})\}\sigma\{\eta(x_{j})\}]$ where $\eta(\cdot)\sim{\textsc{lbp}}(1,b,\mathcal{R})$, denote $\mu(x)=E[\sigma\{\eta(x)\}]$ which is simply $\mu(x)=1/(1+b)$ for any $x$. The mixed moment is
$E\{G_{x_{i}}(B)G_{x_{j}}(B)\}$
=\,\,= $E\left[\sum_{h=1}^{\infty}V_{h}(x_{i})\left\{\prod_{l=1}^{h-1}(1-V_{l}(x_{i}))\right\}\delta_{\theta_{h}}(B)\times\sum_{h=1}^{\infty}V_{h}(x_{j})\left\{\prod_{l=1}^{h-1}(1-V_{l}(x_{j}))\right\}\delta_{\theta_{h}}(B)\right]$
=\,\,= $E\left[\sum_{h=1}^{\infty}V_{h}(x_{i})V_{h}(x_{j})\prod_{l=1}^{h-1}(1-V_{h}(x_{i}))(1-V_{h}(x_{j}))\delta_{\theta_{h}}(B)\right]+$
$E\left[\sum_{h=1}^{\infty}\sum_{l=1}^{h-1}V_{h}(x_{i})\prod_{r=1}^{l-1}\left\{(1-V_{r}(x_{i}))(1-V_{r}(x_{j}))\right\}\{V_{l}(x_{j})-V_{l}(x_{i})V_{l}(x_{j})\}\right.$
×$\left.\times\prod_{s={l+1}}^{h-1}\{1-V_{s}(x_{i})\}\delta_{\theta_{h}}(B)\delta_{\theta_{l}}(B)\right]+$
$E\left[\sum_{h=1}^{\infty}\sum_{l=h+1}^{\infty}V_{l}(x_{j})\prod_{r=1}^{h-1}\left\{(1-V_{r}(x_{i}))(1-V_{r}(x_{j}))\right\}\{V_{h}(x_{i})-V_{h}(x_{i})V_{h}(x_{j})\}\right.$
×$\left.\times\prod_{s={h+1}}^{l-1}\{1-V_{s}(x_{j})\}\delta_{\theta_{h}}(B)\delta_{\theta_{l}}(B)\right]$
=\,\,= $\sum_{h=1}^{\infty}\mu(x_{i},x_{j})\{1-\mu(x_{i})-\mu(x_{j})+\mu(x_{i},x_{j})\}^{h-1}G_{0}(B)+$
$(\mu(x_{j})-\mu(x_{i},x_{j}))\sum_{l=1}^{\infty}\{1-\mu(x_{i})-\mu(x_{j})+\mu(x_{i},x_{j})\}^{l-1}G_{0}(B)^{2}+$
$(\mu(x_{i})-\mu(x_{i},x_{j}))\sum_{h=1}^{\infty}\{1-\mu(x_{i})-\mu(x_{j})+\mu(x_{i},x_{j})\}^{h-1}G_{0}(B)^{2}$
=\,\,= $\frac{\mu(x_{i},x_{j})(G_{0}(B)-G_{0}(B)^{2})}{\mu(x_{i})+\mu(x_{j})-\mu(x_{i},x_{j})}-G_{0}(B)^{2}$
which implies, using $\mu(x)=1/(1+b)$ for any $x$,
$\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}$ $=\frac{\mu(x_{i},x_{j})(G_{0}(B)-G_{0}(B)^{2})}{(\mu(x_{i})+\mu(x_{j})-\mu(x_{i},x_{j}))\times(G_{0}(B)(1-G_{0}(B))/(1+b))}$
$=\frac{(1+b)\mu(x_{i},x_{j})}{2/(1+b)-\mu(x_{i},x_{j})}=\frac{(1+b)^{2}}{2/\mu(x_{i},x_{j})-(1+b)}$
and this result not only holds for single-atoms logistic-beta dependent Dirichlet process, but also for general single-atoms dependent Dirichlet processes with $\mu(x_{i},x_{j})=E\{V(x_{i})V(x_{j})\}$ where $V(x)$ is a distribution of stick-breaking ratios so that $V_{h}(x)$ are independent copies across all $h\in\mathds{N}$. ∎
Next, we present the prior correlation structure under other various dependent Dirichlet process models. In addition to the logistic-beta dependent Dirichlet process (M1), we provide a derivation of correlation lower bounds under three different dependent Dirichlet process models: (M2) squared AR(1) process-based dependent Dirichlet process (DeYoreo and Kottas,, 2018), (M3) time-series dependent Dirichlet process (Nieto-Barajas et al.,, 2012), and (M4) copula-based dependent Dirichlet process. Similar to the logistic-beta dependent Dirichlet process, these models induce dependence through a stochastic process $V(\cdot)$ on stick-breaking ratios that have beta marginals but with different constructions.
First, from DeYoreo and Kottas, (2018, Appendix A.1.), the quantity $\operatorname*{corr}\{V(x_{i}),V(x_{j})\}$ under the squared ar(1) process-based dependent Dirichlet process has the greatest lower bound of $b^{1/2}(b+1)(b+2)^{1/2}-b(b+2)$ when autoregressive parameter converges to 0. This implies $E\{V(x_{i})V(x_{j})\}$ has the infimum of $\{b^{3/2}+(1-b)(b+2)^{1/2}\}/\{(b+1)(b+2)^{1/2}\}$, using $E\{V(x)\}=1/(b+1)$ and $\operatorname*{var}\{V(x)\}=b/((b+1)^{2}(b+2))$ for any $x$ under beta distributed $V(x)$. This leads to the greatest lower bound of correlation between random probability measures $(1+b)/([2(b+2)^{1/2}/\{b^{3/2}+(1-b)(b+2)^{1/2}\}]-1)$, which converges to $0.75$ as $b\to\infty$.
Next, the infimum of $\operatorname*{corr}\{V(x_{i}),V(x_{j})\}$ under the time-series dependent Dirichlet process (Nieto-Barajas et al.,, 2012) is when $V(x_{i})$ and $V(x_{j})$ are independent so that $\inf\operatorname*{corr}\{V(x_{i}),V(x_{j})\}=0$. Thus, $E\{V(x_{i})V(x_{j})\}$ has the infimum of $1/(b+1)^{2}$, which leads to the infimum of correlation between random probability measures $\inf\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}=(1+b)/(1+2b)$. This fact is also noted in Nieto-Barajas et al., (2012) as a special case when stick-breaking ratios are independent, and the same fact also applies to the dependent Dirichlet process proposed by Taddy, (2010) when the autoregressive parameter converges to 0.
Finally, we analyze a copula-based single-atoms dependent Dirichlet process, with beta stick-breaking ratios constructed through a copula that induces a stochastic process with $\mathrm{Beta}(1,b)$ marginals. Finding infimum of $\operatorname*{corr}\{G_{x_{i}}(B),G_{x_{j}}(B)\}$ is equivalent to finding the infimum of correlation between stick-breaking ratios $\operatorname*{corr}\{V(x_{i}),V(x_{j})\}$, which is achieved at the Fréchet lower bound copula (Embrechts et al.,, 2002). In other words, minimum correlation is attained when $V(x_{i})$ and $V(x_{j})$ satisfy $V(x_{j})=F_{B}^{-1}(1-F_{B}(V(x_{i})))$ (counter-monotonic), where $F_{B}$ is the c.d.f. of $\mathrm{Beta}(1,b)$.
Although the analytic form of $\inf\operatorname*{corr}\{V(x_{i}),V(x_{j})\}$ under common atoms logistic-beta dependent Dirichlet process and copula-based dependent Dirichlet process are not readily available due to the nonlinear transformations involved, those can be estimated with Monte Carlo simulation where we used 5,000 simulated datasets of size 1,000 to calculate the correlation estimates.
Figure B.1: Illustration of different correlation lower bounds under four different single-atoms dependent Dirichlet process. (Left) Infimum of correlation between random probability measures under models M1 (solid, black), M2 (dashed, red), M3 (dotted, purple), and M4 (dot-dashed, blue). (Right) Scatterplot of bivariate betas with $\mathrm{Beta}(1,2)$ marginals, corresponding to circles of the left panel.
Figure B.1 illustrates different correlation lower bounds under four different single-atoms dependent Dirichlet processes with concentration parameter $b=2$. The difference in the induced dependence structure is clearly seen from scatterplots, where 1,000 random samples are drawn from a corresponding bivariate beta distribution that induces minimum correlation from models M1–M4. The Fréchet lower bound clearly shows counter-monotonicity between two beta random variables, which induces minimum possible correlation. Transformation of logistic-beta also can induce a negative correlation between beta random variables, while others cannot. In summary, logistic-beta dependent Dirichlet processes can capture a wide range of dependence when $b$ is small to moderate.
The following theorem describes the full-week support property of logistic-beta dependent Dirichlet processes.
Theorem B.1.
Consider a logistic-beta dependent Dirichlet process (10) with an atom process $\{\theta(x):x\in\mathscr{X}\}$ with support $\Theta$ that can be represented with a collection of copulas with positive density w.r.t. Lebesgue measure. Then, the logistic-beta dependent Dirichlet process has full weak support, i.e. the topological support of the logistic-beta dependent Dirichlet process coincides with the space of collections of all probability measures with support $\Theta$ indexed by $x$.
Proof.
Consider the single-atoms logistic-beta dependent Dirichlet process, so that atoms $\theta_{h}$ are i.i.d. samples from the base measure $G_{0}$. Let $\eta(\cdot)\sim{\textsc{lbp}}(1,b,\mathcal{R})$ and $F_{B}$ be the c.d.f. of $\mathrm{Beta}(1,b)$. Consider an $n$-dimensional continuous random variable $U=(F_{B}[\sigma\{\eta(x_{1})\}],\dots,F_{B}[\sigma\{\eta(x_{n})\}])$ where its cumulative distribution $C_{x_{1},\dots,x_{n}}$ is a copula since $U$ has uniform marginals by construction. It is clear that $U$ has a positive density w.r.t. Lebesgue measure since $F_{B}$ and $\sigma$ are all continuous and injective.
Now consider a collection of copula functions $\mathcal{C}=\{C_{x_{1},\dots,x_{n}}:x_{1},\dots,x_{n}\in\mathscr{X},n>1\}$. Since $\sigma$ is the c.d.f. of a standard logistic random variable, the collection of copula functions satisfies conditions of Corollary 1 of Barrientos et al., (2012). Also, the stick-breaking ratios of logistic-beta dependent Dirichlet process $\sigma\{\eta_{h}(x)\}$ are determined by the set of copulas $\mathcal{C}$ and the set of beta marginal distributions with parameters $1$ and $b$. Therefore, the full weak support property of the logistic-beta dependent Dirichlet process follows from Theorem 3 of Barrientos et al., (2012). For a general logistic-beta dependent Dirichlet process where atoms and weights both depend on $x$, the full weak support condition is satisfied as long as the atom process can be represented with a collection of copulas with positive density w.r.t. Lebesgue measure.
Appendix C Additional posterior computation strategies
C.1 Correlation kernels with low-rank structures
Thanks to the normal variance-mean mixture construction of the logistic-beta process, many existing scalable Gaussian process methods that preserve marginal variances can be easily applied to the logistic-beta process. When the dimension of normalized feature map $q$ is less than data size $n$, adopting normalized feature map kernel offers dimension reduction using basis functions (Cressie et al.,, 2022), which is apparent from two representations (6)–(7) in Proposition 3. Specifically, denoting $\Phi=\{\phi_{k}(x_{i})\}_{ik}$ be an $n\times q$ basis matrix with $i$th row $\phi(x_{i})$, the Step 3 of Algorithm 9 becomes sampling $\gamma$ from $q$-dimensional multivariate normal instead of $n$-dimensional multivariate normal,
$\gamma\mid-\sim N_{q}\left((I_{q}+\lambda\Phi^{\top}\Omega\Phi)^{-1}\lambda^{1/2}\Phi^{\top}\{(z_{1:n}-0.51_{n})-0.5\lambda(a-b)\omega\},(I_{q}+\lambda\Phi^{\top}\Omega\Phi)^{-1}\right)$
and set $\eta=0.5\lambda(a-b)1_{n}+\lambda^{1/2}\Phi\gamma$ using representation (7).
Another example of correlation kernel with low-rank structures is the modified predictive process (Finley et al.,, 2009). From a parent kernel $\mathcal{R}$ with $q$ knot locations $u_{1},\dots,u_{q}\in\mathscr{X}$, a new correlation kernel $\tilde{\mathcal{R}}$ with a low-rank structure can be defined as
~$\tilde{\mathcal{R}}(x,x^{\prime})=r(x)^{\mathrm{\scriptscriptstyle T}}R_{uu}^{-1}r(x^{\prime})+\mathds{1}(x=x^{\prime})(1-r(x)^{\mathrm{\scriptscriptstyle T}}R_{uu}^{-1}r(x^{\prime}))$ (14)
where $r(x)=[\mathcal{R}(x,u_{k})]_{k=1}^{q}$ is a length $q$ vector as a function of $x$ and $R_{uu}=[\mathcal{R}(u_{k},u_{k^{\prime}})]_{k,k^{\prime}=1}^{q}$ is a fixed $q\times q$ knot correlation matrix; see also Quiñonero-Candela and Rasmussen, (2005). Thanks to the Woodbury matrix identity and matrix determinant lemma, posterior computation with a low-rank correlation kernel can be carried out in a highly efficient manner. This includes fast evaluation of $n$-dimensional multivariate normal density in Step 2 of Algorithm 9, as the covariance matrix has a “low-rank plus diagonal” structure. In addition to the modified predictive process (14), other options include full-scale approximation (Sang and Huang,, 2012) or multi-resolution approximation (Katzfuss,, 2017) among many others.
Since marginal variances of a new kernel must remain 1 to be a valid correlation kernel, the predictive process (Banerjee et al.,, 2008) or Vecchia approximations (Vecchia,, 1988) cannot be directly applied without suitable adjustments. Exploring other options for scalable logistic-beta processes is an interesting future direction.
C.2 Particle Gibbs sampler
In Step 2 of Algorithm 9, it is possible to use a particle Gibbs sampler (Andrieu et al.,, 2010) to potentially improve the mixing behaviour of $\lambda$. The adaptive Pólya proposal scheme can be used to choose the particle proposal. In Step 2 of Algorithm 9, instead of drawing a single candidate, the particle Gibbs sampler first draws $N$ candidates $\lambda^{(1)},\dots,\lambda^{(N)}\sim\mathrm{Polya}(a^{\prime},b^{\prime})$ and sets the current state as the last candidate $\lambda^{(N+1)}=\lambda$, where $(a^{\prime},b^{\prime})$ are chosen based on adaptive Pólya proposal scheme. Then, unnormalized importance weights $w_{k}=\exp\{-\lambda^{(k)}(ab-a^{\prime}b^{\prime})/2\}g(\lambda^{(k)})$ for $k=1,\dots,N+1$ are calculated, and $\lambda$ is sampled among $N+1$ candidates according to probabilities proportional to $w_{k}$’s. This particle Gibbs sampler also avoids the evaluation of the Pólya density and potentially leads to better mixing with large $N$ but at an increased cost compared to Metropolis-Hastings.
C.3 Challenges for developing rejection samplers involving Pólya densities
Motivated by the broad success of Pólya-Gamma data augmentation algorithms in posterior computation for logistic-type models (Polson et al.,, 2013), we discuss similarities and differences between Pólya and Pólya-Gamma random variables, particularly in the context of rejection samplers. The Pólya-Gamma random variable, denoted as $\omega\sim\textsc{pg}(v,0)$, is defined as an infinite convolution of gamma random variables $(2\pi^{2})^{-1}\sum_{k=0}^{\infty}g_{k}/(k+1/2)^{2},g_{k}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Ga}(v,1)$, and its two-parameter version $\textsc{pg}(v,c)$ arises from an exponential tilting of the $\textsc{pg}(v,0)$ density ${\pi_{\textsc{pg}}}(\omega;v,c)\propto{\pi_{\textsc{pg}}}(\omega;v,0)\exp(-c^{2}\omega/2)$. Compared to Pólya random variables, there is an intersection between Pólya-Gamma where $\mathrm{Polya}(a=1/2,b=1/2)$ is equal in distribution to $\textsc{pg}(v=1,c=0)$ scaled by $4\pi^{2}$, but they generally don’t have overlap for other parameter choices.
The Pólya-Gamma data augmentation scheme facilitates conditionally conjugate updating of the logistic-beta latent parameter $\eta$ in Algorithm 9 and 2, and the development of an efficient rejection sampler for $\textsc{pg}(v,c)$ played an important role in its success in many statistical models. Similarly, the rejection sampler for Kolmogorov–Smirnov outlined in Devroye, (1986, §5.6) has been utilized in the data augmentation scheme of Holmes and Held, (2006). Building rejection samplers for $\textsc{pg}(1,c)$ and Kolmogorov–Smirnov crucially rely on two different alternating sum representations of Jacobi theta functions (Devroye,, 2009), which are essential to meet the monotonicity condition of the rejection sampler based on alternating series (Devroye,, 1986). To the best of our knowledge, we are only aware of such dual representations for two highly special cases of $\mathrm{Polya}(1/2,1/2)$ and $\mathrm{Polya}(1,1)$ that correspond to Pólya-Gamma and Kolmogorov–Smirnov random variables, respectively, but not for general choices of $(a,b)$ of Pólya distribution; see Table 1 of Biane et al., (2001) and Theorems 6.1 and 6.3 of Salminen and Vignat, (2023). Another major hurdle in developing a rejection sampler is that the multivariate normal density in expression (9) does not have a simple form as a function of $\lambda$, in contrast to a simple exponential tilting in the case of Polson et al., (2013). Solving the above problems is an interesting direction for future research.
Appendix D Details of simulation analysis and application study
D.1 Simulation studies
We first provide a detailed description of the quantities we used to evaluate the simulation results. The effective sample size is calculated with an autoregressive approach implemented in R package coda (Plummer et al.,, 2006). Similarly, the multivariate effective sample size is calculated with R package mcmcse (Flegal et al.,, 2017). The root mean squared error and mean absolute error are calculated based on a posterior mean estimate. The mean continuous ranked probability score is calculated by $n^{-1}\sum_{i=1}^{n}S(\hat{F}_{\theta_{i}},\theta_{i}^{*})$, where $S(F,\theta)$ is a continuous ranked probability score (lower the better) and $\hat{F}_{\theta_{i}}$ is an empirical distribution of $\theta_{i}$ obtained from posterior samples of $\theta_{i}$. All algorithms are executed on an Intel(R) Xeon(R) Gold 6132 CPU with a 96GB memory environment. The algorithm running times are defined as the difference between wall-clock time (in seconds) from the start to the finish of Markov chain Monte Carlo algorithm, comprising a total of 2,000 iterations. This excludes pre-computation time, such as the time required to compile Stan programs.
The range parameter of Matérn kernel $\varrho$ is assumed to be unknown, and those are learned from data when fitting both latent logistic-beta process and Gaussian copula models. For the latent logistic-beta process model, we introduce a discrete uniform prior distribution on the set $\{0.01,0.02,\dots,0.5\}$ (with probability 0.02 each), and for the Gaussian copula model, we use continuous uniform prior $\varrho\sim\mathrm{Unif}(0.01,0.5)$ as Stan can only handle continuous priors. The smoothness parameter is fixed as $\nu=1.5$ for both models and the distance calculations $\|x_{i}-x_{j}\|$ in Matérn kernels are based on Euclidean distance.
Table D.1: Simulation results under various latent logistic-beta process inference algorithm settings with latent logistic-beta process data generation scenarios. Entries shown are average over 100 replicates, with Monte Carlo standard error shown in parentheses. LBP, logistic-beta process; M-H, Metropolis-Hastings; ESS, effective sample size of $\lambda$; ESS/sec, ESS divided by running time in seconds.
Data generation Algorithm (blocked) ESS ESS/sec Acceptance rate (%)
Latent LBP, $\varrho=0.1$ Indep. M-H Adapted 245.08 (12.86) 3.35 (0.18) 54.28 (1.03)
Non-adapted 196.35 (11.28) 2.69 (0.15) 49.39 (1.39)
Particle Gibbs Adapted 358.10 (18.57) 2.33 (0.12) 84.50 (0.42)
Non-adapted 352.73 (18.83) 2.34 (0.13) 81.35 (1.02)
Latent LBP, $\varrho=0.2$ Indep. M-H Adapted 257.01 (16.32) 2.89 (0.18) 62.26 (1.12)
Non-Adapted 247.83 (16.53) 2.82 (0.19) 57.60 (1.47)
Particle Gibbs Adapted 360.48 (21.29) 1.96 (0.12) 86.11 (0.40)
Non-adapted 366.55 (24.75) 1.99 (0.13) 84.08 (0.64)
Latent LBP, $\varrho=0.4$ Indep. M-H Adapted 368.26 (17.54) 4.12 (0.20) 66.12 (0.98)
Non-adapted 328.72 (19.00) 3.67 (0.21) 61.45 (1.32)
Particle Gibbs Adapted 468.76 (25.98) 2.52 (0.14) 87.09 (0.29)
Non-adapted 445.09 (25.04) 2.40 (0.14) 85.18 (0.58)
We first present additional simulation results in Table D.1, similar to those in Table 1, to analyze and compare further computational strategies related to the Pólya mixing parameter. Instead of independent Metropolis-Hastings, we employ a particle Gibbs sampler for sampling $\lambda$ with $N=10$ particles, combined with a blocked Gibbs sampler. Table D.1 reports the mixing behaviour and sampling efficiency of algorithms with a particle Gibbs sampler, depending on the utilization of an adaptive Pólya proposal scheme and three different strengths of dependence in data generation scenarios. Here, the acceptance rate for the particle Gibbs sampler is defined as the number of iterations when one of the new $N$ candidates is accepted as a new sample, divided by the total number of iterations.
Table D.1 shows that employing a particle Gibbs sampler for sampling $\lambda$ improves the effective sample size of $\lambda$, but at the cost of decreased sampling efficiency measured by effective sample size divided by seconds. Utilizing the adaptive Pólya proposal with a particle Gibbs sampler shows less significant impact compared to the adaptive Pólya proposal with independent Metropolis-Hastings, as both aim to improve the mixing behaviour of $\lambda$. These results suggest that the particle Gibbs sampler could enhance the mixing behaviour, which is particularly beneficial when the acceptance rate from independent Metropolis-Hastings is low. While this approach may improve mixing, it does not guarantee enhanced sampling efficiency.
Table D.2: Simulation results comparing latent logistic-beta process and Gaussian copula models with latent logistic-beta process data generation scenarios. Entries shown are averaged over 100 replicates, with Monte Carlo standard errors in parentheses. LBP, logistic-beta process; RMSE, root mean squared error; CRPS, continuous ranked probability score; mESS/sec, multivariate effective sample size of latent success probabilities divided by running time in seconds.
Data generation Model RMSE ×$\times 100$ mean CRPS ×$\times 100$ mESS/sec
training test training test
Latent LBP, $\varrho=0.1$ Latent LBP 11.59 (0.15) 12.17 (0.20) 6.31 (0.10) 6.62 (0.12) 21.27 (0.24)
Gauss. copula 11.66 (0.15) 12.19 (0.20) 6.35 (0.09) 6.65 (0.12) 0.48 (0.01)
Latent LBP, $\varrho=0.2$ Latent LBP 8.54 (0.18) 8.73 (0.19) 4.67 (0.11) 4.77 (0.11) 17.87 (0.16)
Gauss. copula 8.59 (0.17) 8.76 (0.18) 4.70 (0.10) 4.79 (0.11) 0.44 (0.01)
Latent LBP, $\varrho=0.4$ Latent LBP 6.12 (0.19) 6.16 (0.19) 3.41 (0.11) 3.43 (0.11) 17.48 (0.16)
Gauss. copula 6.15 (0.17) 6.19 (0.18) 3.43 (0.10) 3.44 (0.11) 0.47 (0.01)
Table D.2 provides simulation results that compare latent logistic-beta process and Gaussian copula models that are similar to Table 2 but with data generation based on the latent logistic-beta process model. Compared to Table 2, where the Gaussian copula model achieved slightly better predictive performance, Table D.2 shows that the latent logistic-beta process model can achieve slightly better predictive performance of of latent success probabilities when the model is correctly specified. In terms of sampling efficiency, the latent logistic-beta process clearly outperforms the Gaussian copula model.
Table D.3: Additional simulation results comparing latent logistic-beta process and Gaussian copula models. Entries shown are average over 100 replicates, with Monte Carlo standard error in parentheses. LBP, logistic-beta process; MAE, mean absolute error of inferred & predicted success probabilities on training & test data; mean ESS/sec, Univariate effective sample sizes of latent success probabilities divided by running time in seconds, averaged across training dataset.
Data generation Model MAE ×$\times 100$ mean ESS/sec time (sec)
training test
Gaussian copula, $\varrho=0.1$ Latent LBP 9.34 (0.13) 9.67 (0.15) 9.44 (0.16) 73.44 (0.48)
Gaussian copula 9.24 (0.12) 9.60 (0.14) 0.49 (0.01) 2135.52 (44.88)
Gaussian copula, $\varrho=0.2$ Latent LBP 6.79 (0.14) 6.90 (0.14) 7.36 (0.15) 89.86 (0.38)
Gaussian copula 6.74 (0.14) 6.86 (0.15) 0.41 (0.01) 2882.76 (255.82)
Gaussian copula, $\varrho=0.4$ Latent LBP 4.82 (0.14) 4.84 (0.14) 7.45 (0.21) 89.98 (0.34)
Gaussian copula 4.80 (0.14) 4.83 (0.14) 0.47 (0.01) 2320.49 (84.65)
Latent LBP, $\varrho=0.1$ Latent LBP 9.03 (0.14) 9.54 (0.17) 9.31 (0.20) 73.50 (0.46)
Gauss. copula 9.10 (0.13) 9.58 (0.17) 0.48 (0.01) 2182.61 (47.37)
Latent LBP, $\varrho=0.2$ Latent LBP 6.64 (0.15) 6.78 (0.16) 7.36 (0.18) 88.65 (0.48)
Gauss. copula 6.67 (0.15) 6.81 (0.15) 0.45 (0.01) 2356.76 (50.61)
Latent LBP, $\varrho=0.4$ Latent LBP 4.84 (0.15) 4.86 (0.16) 7.53 (0.21) 89.91 (0.46)
Gauss. copula 4.85 (0.14) 4.88 (0.15) 0.48 (0.01) 2305.82 (85.17)
We provide additional summaries in Table D.3 that compare latent logistic-beta process and Gaussian copula models. Similar to Table 2, Table D.3 shows that the in-sample and predictive performance of the latent logistic-beta process model in terms of mean absolute error is comparable to the Gaussian copula model, showing the flexibility of logistic-beta process despite model misspecification. When data are generated from the latent logistic-beta process model, the in-sample and predictive performance of the latent logistic-beta process model is slightly better than the Gaussian copula model. We also present sampling efficiency in terms of the average of univariate effective sample size and algorithm running times. Logistic-beta process models have a clear computational advantage, likely due to the conditionally conjugate updating scheme with an efficient blocked Gibbs sampling algorithm.
D.2 Real data analysis
We divide the collaborative perinatal project data available from the BNPmix package into two groups based on maternal smoking status during pregnancy, resulting in data sizes of $n=1023$ (smoking) and $n=1290$ (nonsmoking), respectively. We centred and scaled the data before the analysis, and we set the upper bound on the number of components as $H=20$ for both the logistic-beta dependent Dirichlet process and the logit stick-breaking process. We considered semiconjugate priors for mixture component parameters in expression (13) as $(\beta_{0h},\beta_{1h})^{\mathrm{\scriptscriptstyle T}}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}N_{2}(0_{2},I_{2}),\tau_{h}\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}\mathrm{Ga}(1,1)$, $h=1,\dots,H$ for both models.
We run Algorithm 2 for logistic-beta dependent Dirichlet process, and Algorithm 1 of Rigon and Durante, (2021) for logit stick-breaking process to carry out the posterior inference, both implemented with R for comparison purposes. We ran a total of 35,000 Markov chain Monte Carlo iterations for both models, and the first 5,000 samples were discarded as burn-in, executed on an Intel(R) Xeon(R) Gold 6132 CPU with a 96GB memory environment. The analysis of trace plots shows satisfactory convergence for both models. The running time of Markov chain Monte Carlo samplers under settings 1, 2, and 3 are approximately 5.4, 7.5, and 10.2 minutes for the logistic-beta dependent Dirichlet process and 4.9, 5.6, and 6.2 minutes for the logit stick-breaking process, respectively, for the analysis of smoking group data. This shows the computational tractability of the logistic-beta dependent Dirichlet process model is similar to the logit stick-breaking process model, and we note that substantial time improvements are possible with lower-level programming languages.
The choice of basis functions for logistic-beta dependent Dirichlet process and logit stick-breaking process models with covariate-dependent weights are similar to Rigon and Durante, (2021). We consider a natural cubic spline basis with $q=6$ degrees of freedom as unnormalized feature maps (including the intercept term so that splines are not restricted to meet at origin), obtained from R package splines with automatically chosen knot locations. Basis functions are normalized to have unit Euclidean norm. The left panel of Figure D.1 presents unnormalized and normalized basis functions on the DDE exposure data for the smoking group.
Figure D.1: Comparison of prior properties based on unnormalized (dashed) and normalized feature map (solid). (Left) Unnormalized and normalized basis functions with natural cubic splines for the smoking group. (Right) Comparison of prior co-clustering probabilities and prior expected number of clusters $E(K_{1023})$ given a DDE level under logit stick-breaking process with prior on basis coefficients $\alpha\sim N_{6}(0_{6},\pi^{2}/3I_{6})$.
Under the logit stick-breaking process with unnormalized basis functions $\varphi(x)$ and prior on coefficients $\alpha\sim N_{q}(0_{q},\sigma_{\alpha}^{2}I_{q})$, we have $\eta(x)=\varphi(x)^{\mathrm{\scriptscriptstyle T}}\alpha\sim N(0,\sigma_{\alpha}^{2}\|\varphi(x)\|_{2}^{2}$), whose variance varies over $x$. The corresponding stick-breaking ratio $\sigma\{\eta(x)\}$ follows a logit-normal distribution $F_{V(x)}$ with parameters 00 and $\sigma_{\alpha}^{2}\|\varphi(x)\|_{2}^{2}$, so that prior co-clustering probability $\mu_{2}(x)/(2\mu_{1}(x)-\mu_{2}(x))$ also varies with $x$ where $\mu_{1}(x)=E(V(x))$ and $\mu_{2}=E(V(x)^{2})$ for $V(x)\sim F_{V(x)}$; see Lemma 1 of Dunson and Park, (2008). This is shown as a dashed line in the top right panel of Figure D.1 when $\sigma^{2}_{\alpha}=\pi^{2}/3$ (setting 2), showing marginal prior properties varies with $x$. This is true following previous default choices (Rigon and Durante,, 2021; Horiguchi et al.,, 2022; Kang and Kottas,, 2022). In contrast, under logit stick-breaking process with normalized basis functions $\phi(x)$ and prior $\alpha\sim N_{q}(0_{q},\sigma_{\alpha}^{2}I)$, we have $\eta(x)=\phi(x)^{\mathrm{\scriptscriptstyle T}}\alpha\sim N_{1}(0,\sigma_{\alpha}^{2})$ that does not depend on $x$. The corresponding prior co-clustering probability $\mu_{2}(x)/(2\mu_{1}(x)-\mu_{2}(x))$ also does not depend on $x$, shown as a solid line of the top right panel of Figure D.1 when $\sigma^{2}_{\alpha}=\pi^{2}/3$ (Setting 2). The same fact also holds for the prior expected number of clusters in $n$ samples, denoted as $E(K_{n})$. The bottom right panel of Figure D.1 shows $E(K_{1023})$ under logit stick-breaking process with unnormalized basis (dashed) and normalized basis (solid) functions, obtained from Monte Carlo simulations.
Based on these prior properties, we use normalized basis functions for both the logistic-beta dependent Dirichlet process and the logit stick-breaking process and set their hyperparameters to match their prior co-clustering probabilities. For setting 1, we set $b=0.2$ for the logistic-beta dependent Dirichlet process and $\sigma_{\alpha}^{2}=10^{2}$ for the logit stick-breaking process, which leads to prior co-clustering probabilities $5/6=0.833$ (logistic-beta dependent Dirichlet process) and $0.854$ (logit stick-breaking process), respectively. For setting 2, we set $b=1$ for the logistic-beta dependent Dirichlet process and $\sigma_{\alpha}^{2}=\pi^{2}/3$ for the logit stick-breaking process, where $\pi^{2}/3$ corresponds to the variance of the standard logistic distribution. This leads to prior co-clustering probabilities corresponding to $1/2=0.5$ (logistic-beta dependent Dirichlet process) and $0.515$ (logit stick-breaking process), respectively. For setting 3, we set $b=2$ for the logistic-beta dependent Dirichlet process, $\sigma_{\alpha}^{2}=0.2^{2}$ for the logit stick-breaking process, which leads to prior co-clustering probabilities corresponding to $1/3=0.333$ (logistic-beta dependent Dirichlet process) and $0.338$ (logit stick-breaking process), respectively.
Figure D.2 presents the estimated preterm birth probabilities given DDE exposures for the nonsmoking group of size $n=1290$. Similar to Fig. 5, the estimated preterm birth probabilities show an overall increasing pattern as the DDE exposure level rises. However, unlike the logistic-beta dependent Dirichlet process, the hyperparameter choice of the logit stick-breaking process model has a substantial impact on uncertainty estimates.
Figure D.2: Estimated probability of preterm birth with 95% credible intervals against DDE exposure level from the logistic-beta dependent Dirichlet process model (top) and the logit stick-breaking process model (bottom) with three different hyperparameter settings. DDE exposure data are represented by small vertical bars.